OpenStudy (anonymous):

cube root -9 (times) cube root -81

OpenStudy (watchmath):

9

OpenStudy (anonymous):

\[\sqrt[3]{-9}*\sqrt[3]{-81}=\sqrt[3]{729}=\sqrt[3]{9^3}=9\]